find the 43rd term of each sequence
19.5, 19.9, 20.3, 20.7,
also
21, 15, 9, 3
thank you for ur help

Answer by checkley79(3341)

find the 43rd term of each sequence.
A=a+(n-1)d
19.5, 19.9, 20.3, 20.7,
19.5+(43-1)*(-.6)=19.5+42*-.6=19.5-25.2-5.7 ANS.
also
21, 15, 9, 3
21+(43-1)*(-6)=21+42*-6=21-252=231 ANS.